Shift report 20000903 08:00-16:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Sun Sep 03 2000 - 16:16:38 EDT

  • Next message: http account: "Shift report 20000904 16:00-24:00"

    ------------------------------------------------------------
    Date: Sun Sep  3 16:16:38 2000, Shift: 08:00-16:00
    Supervisor: ZM, with: 
    Log book#5 pages 85 - 88
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam stored and cogged at 10:10 and shortly died.
       Continous beam from 11:25 to 13:30 when was lost
       due to heavy storm - RHIC pass failure.
    
    2. Run# 2519 - 16K trigg#1 collected. Run# 2521 at
       B=19.2E09 and Y=21.4E09, collected 377K trigger#1
    
    3. MRS moved to 60 deg (D5 - 0A). RD measured TPM2 drift
       velocity.
    
    
    
    
     
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Sun Sep 03 2000 - 16:17:13 EDT